OpenStudy (anonymous):

Please help! The equation x^3+px+q=0 has a repeated root. Prove that 4p^3+27q^2=0.

OpenStudy (raden):

if given a polynomial degree's 3 : ax^3 + bx^2 + cx + d = 0 (a,b,c, and d are contants) then : some formula can to solve it x1+x2+x3 = -b/a x1x2 + x1x3 + x2x3 = c/a x1 x2 x3 = -d/a ur equation : x^3+px+q=0 it means, a=4, b=0, c=p, d=q has a repeated root, it means x1=x2=x3 now, we use the formulas above x1+x2+x3 = -b/a x1+x1+x1= 0 3x1 = 0 x1=0 thus x2=0 and x3=0 subtitute x=0 to the original equation : x^3+px+q=0 0^3+p(0)+q=0 q=0 so, x^3+px = 0 next, use the formula : x1x2 + x1x3 + x2x3 = c/a 3(x1)^2 = p 3(0)^2 = p p=0 to prove 4p^3+27q^2=0 just put p=0 and q=0